A study was conducted to determine whether there exists a relationship between the length of time (in hours) a senior high school student spends reviewing his or her lesson a week before the final examination and his or her grade in statistics and probability. Eight randomly selected senior high school students provided the following data.


x 10 12 24 20 13 13 8 22

y 80 83 94 90 85 84 80 89


Compute the Pearson’s r. Interpret the results.


PLEASE ANSWER MY QUESTION QUICKLY!!

DEADLINE : 05/20/2022 11 : 00 PM




1
Expert's answer
2022-05-17T06:48:31-0400

Pearson`s corellation coefficient:

"r=\frac {n\sum (x_iy_i) -\sum {x_i} \sum {y_i}} {\sqrt{(n\sum {x_i^2} -(\sum {x_i})^2)*(n\sum {y_i^2} -(\sum {y_i})^2)}}"


So, we have"\sum_{i=1}^8 x_iy_i=10*80+12*83+24*94+20*90+13*85+13*84+8*80+22*89=10647"

"\sum_{i=1}^8 {x_i}=122" , "\sum_{i=1}^8 {x_i^2}=2106"

"\sum_{i=1}^8 {y_i}=685" , "\sum_{i=1}^8 {y_i^2}=58827"


So, "r=\frac {8*10647-122*685} {\sqrt{8*2106-122^2} *\sqrt{8*58827-685^2}}=0.972"

The range of the correlation coefficient is from -1 to 1. Our result is 0.972 or 97.2%, which means the variables have a positive correlation (it means that for every positive increase in one variable, there is a positive increase of a fixed proportion in the other).